OpenStudy (anonymous):

Two particles move in the xy-plane. At time t, the position of particle A is given by x(t) = 3 t - 3 and y(t) = 2 t - k, and the position of particle B is given by x(t) = 2 t and y(t) = t^2 - 2t - 1. Find k so that the two particles are certain to collide k=?

OpenStudy (anonymous):

to collide xa=xb and ya=yb 3t-3=2t and 2t-k=t^2-2t-1